A home has a triangular backyard. The second angle of the triangle is 4 more than the first angle. The third angle is 26 more than three times the first angle. Find the angles of the triangular backyard
Answer by rfer(16322)

x+x+4+3x+26=180
5x=180-30
5x=150
x=150/5
x=30
x+4=34
3x+26=116